Kuhn Corporation is considering a new project that will require an initial investment of $20,000,000. It has a target capital structure consisting of 45% debt, 4% preferred stock, and 51% common equity.

Kuhn has noncallable bonds outstanding that mature in five years with a face value of $1,000, an annual coupon rate of 10%, and a market price of $1,050.76. The yield on the company’s current bonds is a good approximation of the yield on any new bonds that it issues.

The company can sell shares of preferred stock that pay an annual dividend of $9.00 at a price of $92.25 per share.

Kuhn Corporation does not have any retained earnings available to finance this project, so the firm will have to issue new common stock to help fund it. Its common stock is currently selling for $22.35 per share, and it is expected to pay a dividend of $2.78 at the end of next year. Flotation costs will represent 3.00% of the funds raised by issuing new common stock. The company is projected to grow at a constant rate of 8.70%, and they face a tax rate of 40%.

Kuhn Company’s WACC for this project will be A. 16.64% B.15.95% C.12.48% D.13.87%
Answer #1

Yield using financial calculator
N=5
FV=1000
PV=-1050.76
PMT=10%*1000=100
CPT I/Y=8.70%

After-tax cost of debt=Yield*(1-tax rate)=8.70%*(1-40%)=5.22%

Cost of preferred stock=Dividend/Price=9/92.25=9.76%

Cost of common stock=Expected Dividend/(Price*(1-flotation cost %))+growth rate=2.78/(22.35*(1-3%))+8.70%=21.52%

WACC=45%*5.22%+4%*9.76%+51%*21.52%=13.87%